a dairy owner mixes 5 gal of cream that is 30% butterfat and 8 gal of milk that is 4% butterfat. find the percent concentration of butterfat in the resulting mixture.
Answer by stanbon(75887)

a dairy owner mixes 5 gal of cream that is 30% butterfat and 8 gal of milk that is 4% butterfat. find the percent concentration of butterfat in the resulting mixture.
------
Equation:
fat + fat = fat
------
0.30*5 + 0.04*8 = x*13
----
1.5 + 0.32 = 13x
-----
13x = 1.82
---
x = 14%
======================
Cheers,
Stan H.
================